OpenStudy (anonymous):

3rd root of 324^11, simplifi ed plox

OpenStudy (asnaseer):

first write 324 as a product of its prime factors, so you get:\[324=2^{2}*3^{4}\]so now you have:\[\sqrt[3]{324^{11}}=\sqrt[3]{(2^{2}*3^{4})^{11}}=\sqrt[3]{2^{22}*3^{44}}=\sqrt[3]{2^{21}*2*3^{42}*3^{2}}=2^{7}*3^{14}*\sqrt[3]{2*3^{2}}\]\[\sqrt[3]{324^{11}}=612220032*2^{1/3}*3^{2/3}\]

OpenStudy (anonymous):

Ooops, the actual question is : Multiply then Simplify: